Calculate the pH of a solution obtained by mixing 100 mlof strong acid solution of pH=3 and 400ml solution of same acid solution of pH=1

Grade:12

1 Answers

vikas askiitian expert
13 years ago

millimomes of (acid)1 = [H+]1*100

millimomes of (acid)2 = [H+]2*400

net millimoles = 100( [H+]1+4[H+]2)

[H+]1=10-3      &    [H+]2 = 10-1

net millimoles = 100( 10-3+4*10-1) = (0.401)*100

concentration of resulting solution = total millimomes/total volume(in ml) = 0.401*100/500 = 8*10-2 = [H+] (resulting sol)

PH =-log[H+]=-log8*10-2 =2-log8 = 1.097